A company had developed a forecast using the Delphi method for January through June.  

Month

Demand

Forecast

January

225

240

February

240

230

March

230

235

April

225

230

May

229

225

June

260

230

Using the data provided what is the Mean Absolute Deviation for the forecast method?

3.17

-3.17

11.5

69

Using the data provided, what is the Mean Squared Error?

215.17

20.5%

9.4%

12%

Using the data provided what is the MAPD?

5.7%

2.1%

4.9%

3.6%
Answer #1

Solution:

Given:

A company had developed a forecast using the Delphi method for January through June.  

Month

Demand

Forecast

January

225

240

February

240

230

March

230

235

April

225

230

May

229

225

June

260

230

Part a) We have to find the Mean Absolute Deviation for the forecast method

MAD D -F

Thus we need to make following table:

Dt Ft | Dt - Ft |
225 240 15
240 230 10
230 235 5
225 230 5
229 225 4
260 230 30
\sum \left | D_{t}-F_{t} \right | = 69

Thus

MAD = \frac{\sum \left | D_{t}-F_{t} \right |}{n}

MAD = \frac{69}{6}

MAD = 11.5

Part b) Find the Mean Squared Error

MSE = \frac{\sum \left ( D_{t}-F_{t} \right )^{2}}{n}

Thus we need to maje following table:

Dt Ft Dt - Ft (Dt - Ft )^2
225 240 -15 225
240 230 10 100
230 235 -5 25
225 230 -5 25
229 225 4 16
260 230 30 900
\sum \left ( D_{t}-F_{t} \right )^{2}= 1291

MSE = \frac{\sum \left ( D_{t}-F_{t} \right )^{2}}{n}

MSE = \frac{1291 }{6}

MSE = 215.17

Part c) Find the MAPD

MAPD = \frac{\sum \left | D_{t}-F_{t} \right |}{\sum D_{t}}

Thus

Dt Ft | Dt - Ft |
225 240 15
240 230 10
230 235 5
225 230 5
229 225 4
260 230 30
\sum D_{t}=1409 \sum | D_{t} - F_{t}|=69

Thus

MAPD = \frac{\sum \left | D_{t}-F_{t} \right |}{\sum D_{t}}

MAPD = \frac{69}{1409}

MAPD = 4.9 \%